OpenStudy (anonymous):

The base and height of a triangle are tripled. What is the effect on its area and perimeter?

OpenStudy (anonymous):

original area is A=(1/2)bh the new area is A1 = (1/2)(3b)(3h) A1 = 9*(1/2)(b)(h) A1 = 9*A the new area is now 9 times as big as the original.